# Why is Grain Growth Not Curvature Flow?

### v1

@@ II.3 Failure of mean curvature flow
The von Neumann-Mullins relation (Eq. (1)) stems from the mean curvature flow assumption, $v=M\gamma\kappa$, i.e., GB velocity is proportional to the mean GB curvature. However, statistical analysis of experimental measurements on nickel [15] and iron [16] showed that GB velocity is not (or only weakly) correlated with mean curvature (data reproduced in Figs. 4a and 4b). Similar conclusions were drawn from MD simulations of polycrystalline aluminum (Al) [17]; see Fig. 4c. The small colored dots in these figures are individual measurements of the velocity and (weighted) mean curvature of different GBs; the color of each dot indicates the density of dots (measurements) in the local vicinity of each dot (redder dots indicate there are many GBs with similar velocities and curvatures. The open black circles refer to the average GB velocity of all GBs with the same (weighted) mean curvature (binned data).
@@
Clearly, any relationship between the GB velocity and mean curvature is very weak; certainly $v=M\gamma\kappa$ is inconsistent with these individual GB data. This is shown by noting the very small correlation between the average GB velocity and mean GB curvature (see the open circles); the coefficient of determination (i.e., the correlation) $COD$ for the experiments and simulations are all less than $0.35$. Since the GB velocity and GB mean curvature are only weakly correlated, the fundamental assumption that led to the von Neumann-Mullins relation (or its extensions to higher dimensions or inclusion of anisotropic GB energy) is invalid. Note that while the GB velocity is not proportional to GB curvature, capillarity (reduction in GB energy) remains the driving force for normal grain growth; curvature flow is necessarily modified by shear coupling.
@@
Figure 4: Failure of mean curvature flow. Scatter plots (small points) of GB velocity vs. mean curvature for individual GBs in polycrystalline (a) Ni (experiment), (b) $\alpha$-Fe (experiment) and (c) Al (MD simulation). The color of each of the small points indicates the density of points (measurements) in the local vicinity of each point (redder points indicate there are many GBs with similar velocities and curvatures. The open circles represent GB velocities averaged over GB segments with (approximately) the same mean curvature. (d) Illustration of GB migration mediated by disconnection motion in opposite directions. (e,f,g) Same type of plots (a)-(c) for PF grain growth simulations with shear coupling for polycrystals with shear modulus (e) $\mu=75$ GPa, (f) $\mu=25$ GPa and (g) with no shear coupling. $\mu=0$. (h)-(i) Same as (g) but with (h) anisotropic GB energy and (i) with both anisotropic GB energy and triple junction drag (see the text). (a)-(c) are reproduced with permission from [15, 16, 17] (Copyright 2021 AAAS, 2023 Elsevier, 2024 Elsevier).
@@
Why is the GB velocity not (only weakly) correlated with GB mean curvature? The simulation results above strongly suggest that this is related to shear coupling. As noted (see Fig. 2a), GB migration and shear coupling are mediated by the motion of disconnections. A more appropriate equation of motion for GBs, that accounts for shear coupling in addition to curvature effects, can be expressed as [22, 23]
@@
Eq. (2) implies that GBs can migrate either toward or against its curvature, due to the existence of internal stresses. Consider the situation in the schematic in Fig. 4d where the capillarity force drives the GB segment downward (toward its center of curvature) via the motion of the two vertical disconnections towards each other. However, the internal stress induced by one disconnection leads to elastic interactions in terms of a Peach-Koehler force on the Burgers vector of the other disconnection. If the total Peach-Koehler force acting on the disconnections drives them toward each other, the GB segment will migrate in the same direction as when driven by curvature albeit faster. On the other hand, if the Peach-Koehler force drives these two disconnections apart, the resultant driving force is opposite to that of mean curvature flow; for sufficiently large Peach-Koehler forces the GB segment will migrate in the direction opposite to that in curvature flow. In any case, this elastic effect weakens the correlation between GB velocity and curvature.
@@
We now demonstrate that the main cause for the weak correlation between GB velocity and curvature is, indeed, the generation of internal stresses associated with GB shear coupling. We analyze the GB velocity vs. GB mean curvature distributions obtained by our PF grain growth simulations with and without GB shear coupling, $c.f.$ Figs. 4e-g, where we employed the Bhattacharya et al. [15] approach to extract GB velocity and mean curvature during grain growth. When grain growth is mean curvature flow without GB shear coupling (Fig. 4g), a clear linear distribution and very strong correlation ($COD\approx 1$) between GB velocity and mean curvature are found; this is unsurprising (note that while this is mean curvature flow, each data point corresponds to an average over an entire GB (rather than points on a GB); this explains the sparse deviation from the exact, isotropic, von Neumann-Mullins law). Figures 4e and 4f show what happens when GB shear coupling is included; the correlation between curvature and velocity is not only lost, but the resultant distribution of curvature-velocity scatter appears very similar to the experimental and MD results (c.f., Figs. 4a-c and e,f). The strong similarity is observed both for individual GBs and for the results averaged over GBs with the same curvature, whilst the $COD$ from the PF simulations are similar to those from the experimental data. Shear coupling generates Peach-Koehler forces which are mediated by elastic fields, hence small shear modulus $\mu$ would imply decreased shear coupling effects. We performed PF simulations with small ($\mu=25$ GPa) and large ($\mu=75$ GPa) shear moduli (c.f. , Figs. 4e and f) and, indeed, find that reducing the shear modulus increases the correlation between GB velocity and curvature. This is consistent with the fact that Fe and Ni have high modulus but low $COD$ and Al has low modulus and higher $COD$.
@@
The results presented above provide strong, clear evidence that the (experimental and MD simulations) observations of weak correlation between GB velocity and curvature is a direct consequence of GB shear coupling. Other possible reasons for this lack of correlation have been proposed: (i) GB replacement as a result of anisotropic GB energy [36] and (ii) triple junction drag effect [16]. To examine the effect of anisotropic GB energy, we conducted a series of PF simulations that include misorientation and inclination-dependent anisotropic GB energies without GB shear coupling (using the GB energy function of Al from [37]); the resultant GB velocity vs. mean curvature distribution is shown in Fig. 4h. The results are similar to the mean curvature flow simulation results (Fig. 4g), albeit with a slightly weaker curvature-velocity correlation. Hence, GB energy anisotropy does lead to weaker velocity-curvature correlation but still very much higher than that found in either the experiments or MD simulation. Note that GB energy anisotropy indeed leads to the replacement of high-energy GBs with low-energy GBs in our PF simulations (see an example in Fig. S2 in SI), as in the experiments [15].
@@
We also investigated the effects of triple junction drag on the GB velocity-curvature correlation by reducing the TJ mobility (i.e., by lowering the PF kinetic coefficient in the triple junction region by 10$\times$ [38]). As noted by Gottstein and Shvindlerman [39], when grain growth is controlled by TJ drag, the grain growth law is completely different from the von Neumann-Mullins law. In our PF simulations, we find that when the GB mean curvature is low, the curvature-velocity correlation is high, but when the mean curvature is large, the correlation is notably weaker (see Fig. 4i). Nonetheless, including TJ drag does, indeed, lower the curvature-velocity correlation but the $COD$ remains much larger than that found in the experiments or MD simulations. It is also interesting to note that the individual GB velocity-curvature measurements (small colored points in Fig. 4i) are distributed much differently than those in the experiments or MD simulations (Fig. 4a-c). Hence, we conclude that TJ drag is not the dominant cause of the failure of curvature flow.
While anisotropic GB energy and TJ drag reduce the correlation between GB velocity and curvature, it is clear that shear coupling is the main source of the failure of curvature flow in grain growth.
@@
Figure 5: Model microstructures for benchmarking. (a) Schemetic of the initial [111]-textured idealized microstructure for the reported MD and current PF grain growth simulations in (b)-(e). Microstructure evolution in (b) reported MD simulations at time around 5 ns and (d) current PF simulations at time of 15000. Internal stress field $\sigma_{\text{12}}$ at time (c) 5.5 ns in the MD simulations and (e) 15000 in the PF simulations. (f) Schemetic of the initial [001]-textured idealized microstructure for the MD and PF grain growth simulations in (g)-(j). Microstructure evolution in (g) MD simulations at time around 20 ns and (h) PF simulations at time of 2500. Internal stress field $\sigma_{\text{12}}$ at time (i) 20 ns in the MD simulations and (j) 2500 in the PF simulations. (b), (c) are reproduced with permission from [21] (Copyright 2017 Springer Nature)
III Discussion
@@
We conduct PF simulations on the same idealized microstructure using material parameters determined from bicrystal MD simulations [21, 29] (see the SI); i.e., there are no adjustable physical parameters. Without GB shear coupling, the shrinkage rates of grains A and B are, unsurprisingly, the same (see Figs. S3a,c in SI); i.e., this microstructure evolution is not consistent with the MD simulations results. Next, we performed PF simulations on the same microstructure but where we assign the measured shear coupling factors (from bicrystal MD simulations) to the individual GBs. Detailed comparisons of the PF (Figs. 5d,e) and MD (Figs. 5b,c) microstructure simulations show that the PF simulations with shear coupling accurately reproduce the MD microstructure evolution, internal stress fields and the difference in shrinkage rates of grain A and B.
@@
We also consider a second example of the evolution of this idealized microstructure with different texture (tilt axis along thickness direction) and initial grain orientations, see the schematic initial configuration in Fig. 5f; here too, the square columnar grains A and B share the same orientation. In MD simulations, the shrinkage rates of grains A and B are almost the same, whilst their shapes are elongated along one diagonal and compressed along the other; see Fig. 5g. We also find the different stresses in grains A and B (see Fig. 5h). By assigning the shear coupling factors to inidividual GBs, the microstructure evolution and internal stress field obtained by our PF simulation match very well again with the MD simulation, c.f. Figs. 5i,j and Figs. 5g,h. These results (i) demonstrate that the PF model with shear coupling provides an accurate representation of polycrystalline microstructure evolution and (ii) further demonstrate that the observation that GB velocity is not proportional to GB curvature (in the presence of shear coupling) translates into the failure of the von Neumann-Mullins law Eq. (1) at the microstructural level.
@@
We examined the effects of GB shear coupling on the microstructure evolution through a series of large-scale PF simulations of grain growth in two-dimensional (2D) polycrystals. The PF simulations are based on a bicrystallography-respecting continuum model for disconnection-mediated GB migration (see the SI for a discussion of the physical parameters in the PF simulations). We can now definitively answer the question asked in the title: “why is grain growth not curvature flow?” We showed that GB shear coupling/disconnections is responsible for pronounced deviations from classical curvature-driven grain growth; shear coupling generates internal stress which acts as additional (non-capillarity) driving forces for GB migration. These mechanical driving forces modify the classical GB equation of motion (i.e., mean curvature flow) thereby rendering the classical 2D von Neumann-Mullins model (and its extensions to higher dimensions) unable to describe grain growth. Nonetheless, we emphasize that the ultimate driving force for grain growth remains the reduction in GB energy (capillarity), as in classical models, however inclusion of elasticity/shear coupling is essential in the GB equation of motion in order to describe the kinetic process of grain growth.
IV Methods

### v3

@@ I.3 Failure of mean curvature flow
The von Neumann-Mullins relation (Eq. (1)) stems from the mean curvature flow assumption, $v=M\gamma\kappa$, i.e., GB velocity is proportional to the mean GB curvature. However, statistical analysis of experimental measurements on nickel [18] and iron [19] showed that GB velocity is not (or only weakly) correlated with mean curvature (data reproduced in Figs. 4a and b). Similar conclusions were drawn from MD simulations of polycrystalline aluminum (Al) [20]; see Fig. 4c. The small colored dots in these figures are individual measurements of the velocity and (weighted) mean curvature of different GBs; the color of each dot indicates the density of dots (measurements) in the local vicinity of each dot. Redder dots indicate there are many GBs with similar velocities and curvatures. The open black circles refer to the average GB velocity of all GBs with the same (weighted) mean curvature (binned data).
@@
Clearly, any relationship between the GB velocity and mean curvature is very weak; certainly $v=M\gamma\kappa$ is inconsistent with these individual GB data. This is shown by noting the very small correlation between the average GB velocity and mean GB curvature (see the open circles); the coefficient of determination $COD$ (i.e., the linear correlation, often called $R$ squared) for the experiments and simulations are all less than $0.35$. Since the GB velocity and GB mean curvature are only weakly correlated, the fundamental assumption that led to the von Neumann-Mullins relation (or its extensions to higher dimensions or inclusion of anisotropic GB energy) is invalid. Note that while the GB velocity is not proportional to GB curvature, capillarity (reduction in GB energy) remains the driving force for normal grain growth; curvature flow is necessarily modified by shear coupling.
@@
Figure 4: Failure of mean curvature flow. Scatter plots (small points) of GB velocity vs. mean curvature for individual GBs in polycrystalline (a) Ni (experiment), (b) $\alpha$-Fe (experiment), and (c) Al (MD simulation). The color of each of the small points indicates the density of points (measurements) in the local vicinity of each point (redder points indicate there are many GBs with similar velocities and curvatures. The open circles represent GB velocities averaged over GB segments with (approximately) the same mean curvature. (d) An example of GB migration opposite to its center of mean curvature in grain growth experiments in (b). (e) Illustration of GB migration mediated by disconnection motion in opposite directions. (f) Simulated grain growth microstructure evolution with GB shear coupling ($\mu=25$ GPa). Here, increasing time is indicated by changes in the GB colors from blue to red. Black arrows indicate locations where the GBs migrate in the sense opposite to the curvature. (g,h,i) Same type of plots (a)-(c) for PF grain growth simulations with shear coupling for polycrystals with shear modulus (e) $\mu=75$ GPa, (f) $\mu=25$ GPa and (g) with no shear coupling. $\mu=0$. (j)-(l) Same as (i) but with (j) anisotropic GB energy, (k) anisotropic GB mobility and (l) triple junction drag (see the text). (a)-(d) are reproduced with permission from [18, 19, 20] (Copyright 2021 AAAS, 2023 Elsevier, 2024 Elsevier).
@@
Why is the GB velocity not (only weakly) correlated with GB mean curvature? The simulation results above strongly suggest that this is related to shear coupling. As noted (see Fig. 2a-b), GB migration and shear coupling are mediated by the motion of disconnections. A more appropriate equation of motion for GBs, that accounts for shear coupling in addition to curvature effects, can be expressed as [25, 26]
@@
Eq. (2) implies that GBs can migrate either toward or against its curvature due to the existence of internal stresses. Consider the situation in the schematic in Fig. 4e where the capillarity force drives the GB segment downward (toward its center of curvature) via the motion of the two vertical disconnections towards each other. However, the internal stress induced by one disconnection leads to elastic interactions in terms of a Peach-Koehler force on the Burgers vector of the other disconnection. If the total Peach-Koehler force acting on the disconnections drives them toward each other, the GB segment will migrate in the same direction as when driven by curvature albeit faster. On the other hand, if the Peach-Koehler force drives these two disconnections apart, the resultant driving force is opposite to that of mean curvature flow; for sufficiently large Peach-Koehler forces the GB segment will migrate in the direction opposite to that in curvature flow. In any case, this elastic effect weakens the correlation between GB velocity and curvature. GB migration in the direction opposite to the center of curvature has been observed in grain growth experiments [19, 35]; e.g., see Fig. 4d from Ref. [19]. In our grain growth simulations with GB shear coupling, we also observe GB motion in the direction opposite to the center of the curvature; see the GB network trajectory in Fig. 4f where indicated by the arrows. (Increasing time in this microstructure is indicated by the transition of the GB color from blue to red.) Note that, while anisotropic $\mathbf{M}$ or $\Gamma$ in Eq. (2) (i.e., weighted mean curvature flow) are essential for accurately describing the magnitude of the velocity, they cannot capture these observations of GB migration in the direction opposite to the sense of the curvature, as illustrated in Figs. 4d-e, see also Figs. S4c-f in SI. Without shear coupling, such motion might still occur due to the torque term in $\Gamma$ for high-energy GBs [39]. This effect could then be observed in the early stages of the evolution of microstructures featuring randomly oriented GBs. As experiments typically show faceted GBs with orientations close to minimum-energy ones rather than random, the significant amount of GBs found to move opposite to curvature [19] cannot be explained by this effect only.
@@
We now demonstrate that the main cause for the weak correlation between GB velocity and curvature is, indeed, the generation of internal stresses associated with GB shear coupling. We analyze the GB velocity vs. GB mean curvature distributions obtained by our PF grain growth simulations with and without GB shear coupling, $c.f.$ Figs. 4g-i, where we employed the Bhattacharya et al. [18] approach to extract GB velocity and mean curvature during grain growth. When grain growth is mean curvature flow without GB shear coupling (Fig. 4i), a clear linear distribution and very strong correlation ($COD\approx 1$) between GB velocity and mean curvature are found; this is unsurprising (note that while this is mean curvature flow, each data point corresponds to an average over an entire GB, rather than points on a GB; this explains the sparse deviation from the exact, isotropic, von Neumann-Mullins law). Figures 4g and h show what happens when GB shear coupling is included; the correlation between curvature and velocity is not only lost, but the resultant distribution of curvature-velocity scatter appears very similar to the experimental and MD results ($c.f.$, Figs. 4a-c and g,h). The strong similarity is observed both for individual GBs and for the results averaged over GBs with the same curvature, whilst the $COD$ from the PF simulations are similar to those from the experimental data. Shear coupling generates Peach-Koehler forces which are mediated by elastic fields, hence small shear modulus $\mu$ would imply decreased shear coupling effects. We performed PF simulations with small ($\mu=25$ GPa) and large ($\mu=75$ GPa) shear moduli ($c.f.$ , Figs. 4g and h) and, indeed, found that reducing the shear modulus increases the correlation between GB velocity and curvature. This is consistent with the fact that Fe and Ni have a high modulus but low $COD$, and Al has a low modulus and higher $COD$.
@@
The results presented above provide strong, clear evidence that the (experimental and MD simulations) observations of weak correlation between GB velocity and curvature is a direct consequence of GB shear coupling. Other possible reasons for this lack of correlation have been proposed: (i) GB replacement as a result of anisotropic GB energy [40] or anisotropic GB mobility [41] and (ii) triple junction drag effect [19]. To examine the effect of anisotropic GB energy, we conducted a series of PF simulations that include misorientation and inclination-dependent anisotropic GB energies without GB shear coupling (using the GB energy function of Al from [42]); the resultant GB velocity vs. mean curvature distribution is shown in Fig. 4j. The results are similar to the mean curvature flow simulation results (Fig. 4i), albeit with a slightly weaker curvature-velocity correlation. Hence, GB energy anisotropy does lead to weaker velocity-curvature correlation but still very much higher than that found in either the experiments or MD simulation. Note that GB energy anisotropy indeed leads to the replacement of high-energy GBs with low-energy GBs in our PF simulations (see an example in Fig. S3 in SI), as in the experiments [18]. Figure 4k shows the results of simulations performed with isotropic GB energy but anisotropic GB mobility, where the magnitude of the GB mobility anisotropy is consistent with the experiments of Zhang et al. [41]; see SI for details. The results for the simulations with anisotropic GB mobility is very similar to those for anisotropic GB energy.
@@
We also investigated the effects of triple junction drag on the GB velocity-curvature correlation by reducing the TJ mobility (i.e., by lowering the PF kinetic coefficient in the triple junction region by 10$\times$ [43]). As noted by Gottstein and Shvindlerman [44], when grain growth is controlled by TJ drag, the grain growth law is completely different from the von Neumann-Mullins law. In our PF simulations, we find that when the GB mean curvature is low, the curvature-velocity correlation is high, but when the mean curvature is large, the correlation is notably weaker (see Fig. 4l). Nonetheless, including TJ drag does, indeed, lower the curvature-velocity correlation, but the $COD$ remains much larger than that found in the experiments or MD simulations. It is also interesting to note that the individual GB velocity-curvature measurements (small colored points in Fig. 4l) are distributed much differently than those in the experiments or MD simulations (Figs. 4a-c). Hence, we conclude that TJ drag is not the dominant cause of the failure of curvature flow.
@@
Figure 5: Model microstructures for benchmarking. (a) Schemetic of the initial [111]-textured idealized microstructure for the reported MD and current PF grain growth simulations in (b)-(e). Microstructure evolution in (b) reported MD simulations at time around 10 ns and (d) current PF simulations at time of 15000. Internal stress field $\sigma_{\text{12}}$ at time (c) 5.5 ns in the MD simulations and (e) 7500 in the PF simulations. (f) Schemetic of the initial [001]-textured idealized microstructure for the MD and PF grain growth simulations in (g)-(j). Microstructure evolution in (g) MD simulations at time around 20 ns and (h) PF simulations at time of 2000. Internal stress field $\sigma_{\text{12}}$ at time (i) 20 ns in the MD simulations and (j) 2000 in the PF simulations. (b), (c) are reproduced with permission from [24] (Copyright 2017 Springer Nature)
While anisotropic GB energy, mobility, and TJ drag reduce the correlation between GB velocity and curvature, it is clear that shear coupling is the main source of the failure of curvature flow in grain growth.
II Discussion
@@
We conduct PF simulations on the same idealized microstructure using material parameters determined from bicrystal MD simulations [24, 32] (see the SI); i.e., there are no adjustable physical parameters. Without GB shear coupling, the shrinkage rates of grains A and B are, unsurprisingly, the same (see Figs. S5a,c in SI); i.e., this microstructure evolution is not consistent with the MD simulations results. Next, we performed PF simulations on the same microstructure, but we assigned the measured shear coupling factors (from bicrystal MD simulations) to the individual GBs. Detailed comparisons of the PF (Figs. 5d,e) and MD (Figs. 5b,c) microstructure simulations show that the PF simulations with shear coupling accurately reproduce the MD microstructure evolution, internal stress fields, and the difference in shrinkage rates of grain A and B.
@@
A second example of the evolution of this idealized microstructure with different texture (tilt axis along thickness direction) and initial grain orientations results in analogous conclusions, see the schematic initial configuration in Fig. 5f; here too, the square columnar grains A and B share the same orientation. In MD simulations, the shrinkage rates of grains A and B are almost the same, whilst their shapes are elongated along one diagonal and compressed along the other; see Fig. 5g. We also find the different stresses in grains A and B (see Fig. 5h). By assigning the shear coupling factors to individual GBs, the microstructure evolution and internal stress field obtained by our PF simulation match very well again with the MD simulation, $c.f.$ Figs. 5i,j and Figs. 5g,h. These results (i) demonstrate that the PF model with shear coupling provides an accurate representation of polycrystalline microstructure evolution and (ii) further demonstrate that the observation that GB velocity is not proportional to GB curvature (in the presence of shear coupling) translates into the failure of the von Neumann-Mullins law, Eq. (1), at the microstructural level.
Increasing temperature tends to decrease the magnitude of the GB shear coupling as a result of activation of multiple disconnection modes [32, 46, 23, 47, 29]. Such a weakened shear coupling factor can improve the correlation between $v$ and $\kappa$, similar to the effect by decreasing the shear modulus as shown in Figs. 4g and h (shear moduli also slowly decrease with increasing temperature). In the limit that $T\rightarrow\infty$ (ignoring melting), grain growth can be considered as curvature flow. The magnitude of this effect is material- (and GB-) dependent. Atomistic simulations show that grain growths in Al polycrystal at 0.75 $T_{\rm m}$ [20] and Ni polycrystal at 0.85 $T_{\rm m}$ [24] are still significantly different from curvature flow. Hence, at most temperatures where grain growth experiments and simulations are conducted, GB shear coupling cannot be ignored to predict and understand realistic grain growth.
@@
In summary, we examined the effects of GB shear coupling on the microstructure evolution through a series of large-scale PF simulations of grain growth in two-dimensional (2D) polycrystals. The PF simulations are based on a bicrystallography-respecting continuum model for disconnection-mediated GB migration (see the SI for a discussion of the physical parameters in the PF simulations). The inclusion of shear coupling (disconnection) effects provides a clear explanation of why grain growth is not weighted mean curvature flow (these are captured in the simulations). More specifically, shear coupling generates internal stresses that provide important additional (to capillarity) driving forces for grain boundary migration. These mechanical driving forces modify the classical GB equation of motion (i.e., mean curvature flow and its anisotropic extensions), thereby rendering the classical 2D von Neumann-Mullins model (and its extensions to higher dimensions and anisotropy) unable to describe grain growth. While the effects of anisotropic GB energy and mobility do, indeed, decrease the correlation between GB velocity and weighted mean curvature, they are too weak to account for the observed deviations (both in magnitude and trends) of the experimental and simulation data from weighted mean curvature flow. We emphasize that the ultimate driving force for grain growth remains the reduction in GB energy (capillarity), as in classical models, but the inclusion of elasticity/shear coupling is essential in the GB equation of motion in order to describe the kinetic process of grain growth.
III Methods